Carla Vista, Inc., has four-year bonds outstanding that pay a coupon rate of 8.00 percent and make coupon payments semiannually. If these bonds are currently selling at $920.89. What is the yield to maturity that an investor can expect to earn on these bonds? (Round answer to 1 decimal place, e.g. 15.2%.) Yield to maturity % What is the effective annual yield? (Round answer to 1 decimal place, e.g. 15.2%.) Effective annual yield
Answer #1

Information provided:
Future value= $1,000

Time= 4 years*2= 8 semi-annual periods

Coupon rate= 8%/2= 4% per semi-annual period

Coupon payment= 0.04*1,000= $40

Present value= $920.89

The yield to maturity is calculated by entering the below in a financial aclcualtor:

FV= 1,000

N= 8

PMT= 40

PV= -920.89

Press the CPT key and I/Y to compute the yield to maturity.

The value obtained is 5.24%.

Therefore, the yield to maturity is 5.24% for a semi-annual period and 5.24%*2= 10.47% 10.5% for an annual period.

The effective annual yield is calculated using the below formula:

Effective annual yield= (1+r/n)^n-1

Where r is the interest rate and n is the number of compounding periods in one year.

Effective annual yield = (1+0.1047/2)^2-1

                                   = 1.1074-1

                                    = 0.1074*100

                                     = 10.74%.

Therefore, the effective annual yield is 10.74% 10.7%.